OpenStudy (djbreezy):

Which ordered pair is a solution of -3x + y = -5? (8, -16) (8, -29) (8, -64) (8, 19)

OpenStudy (welshfella):

just plug in the given values of x and y from each option into the equation and see which ones fit.

OpenStudy (luckyness):

Alright so a ordered pair is when you can take out x and y and put in there values, and then the equation = to true. Lets try A: -3x + y = -5 Replace -3*8 + -16 = -5 ~False, so not A. Lets try B: -3x + y = -5 Replace: -3*8 + -29 = -5 ~False, so not B. Lets try C: -3x + y = -5 Replace: -3*8 + -64 = -5 ~False, so not C. Leaving us with D, but lets check: D: -3x + y = -5 Replace: -3*8 + 19 = -5 ~True! Your final answer is D.

OpenStudy (djbreezy):

thx

OpenStudy (luckyness):

Your welcome!